OpenStudy (anonymous):

Find the larger root for 6x^2−5=3

OpenStudy (campbell_st):

add 5 to both sides 6x^2 = 8 divide by 6 x^2 = 4/3 \[x = \pm \sqrt{4/3} = \pm 2/\sqrt{3}\] or \[x = \pm (2\sqrt{3})/3\] the larger is the positive... there are 3 versions of the answer \[x = \sqrt{4/3}\] \[x = 2/\sqrt{3}\] \[x = (2\sqrt{3})/3\]

OpenStudy (campbell_st):

each version has the same numeric value

OpenStudy (anonymous):

i dont know but , i keep getting it wrong :'(